Two numbers differ by 57. Their sum is 185. Find the numbers. I don't understand this whole process. Our teacher wants all the steps. HELP!
Answer by tinbar(133)

Let the two number be x and y.
Then x-y = 57 (1)
and x+y = 185 (2)
From (1) x-y=57, we can re-arrange and get x = 57+y.
Now since the x and y in both (1) and (2) are the same x and y, we can take our x=57+y and substitute that value for the x in (2).
So now our (2) goes from x+y=185 to (57+y)+y=185. Now we can solve for y, 57+2y=185,2y=128, y = 64. So now we can plug in this value for y in either (1) or (2) doesn't matter, to find the value of x. From (1) we get x-64=57, so x =57+64 = 121.
So now let's check, 121-64 = 57 is true, so our first statement is satisfied. I'll leave it to you to check the second statement. If that is also satisfied, then you have your answer!